# MetaML-Pro: Cross-Stage Design Flow Automation for Efficient Deep Learning Acceleration

### v1

@@ 4.6. Design Space Exploration (DSE)
5. Evaluation
5.1. Experimental Setup
5.2. Optimization Strategy using Single $O$-task
@@
VGG7
Baseline
4568
37.2%
345k
20.0%
65k
1.9%
98.2%
VGG7
Quant.
0.01
934
7.6%
313k
18.1%
51k
1.5%
97.5%
VGG7
Quant.
0.05
505
4.1%
318k
18.4%
66k
1.9%
93.9%
JetDNN
Baseline (Que et al., 2023)
638
290%
66k
124%
30k
28%
76.1%
JetDNN
Quant.
0.01
75
34.1%
57k
107%
17k
16%
75.5%
JetDNN
Quant.
0.05
61
27.7%
76k
144%
7.2k
6.8%
71.6%
@@
5.3. Custom Optimization Strategy using Multiple $O$-tasks
@@ 5.6. Bottom-up Flows
5.7. Optimization strategy search - combination and order
@@
5.8. Optimization strategy search - tolerable loss variation
@@ 5.9. DSE Strategies
Model
$\alpha_{s}$, $\alpha_{p}$, $\alpha_{q}$
FPGA
Acc.
Lat.
DSP
LUT
HLS4ML
Jet-DNN (Duarte et al., 2018)
- KU115 75 75
954
FPL’20 (Umuroglu et al., 2020)
LogicNets JSC-M
- VU9P 70.6 NA 0 (0)
14,428
FPL’20 (Umuroglu et al., 2020)
LogicNets JSC-L
- VU9P 71.8 13a 0 (0)
37,931
NMI’21 (Coelho et al., 2021)
Qkeras Q6
- VU9P 74.8 55
124
39,782
NMI’21 (Coelho et al., 2021)
AutoQkeras QE
- VU9P 72.3 55
66
9,149
NMI’21 (Coelho et al., 2021)
AutoQkeras QB
- VU9P 71.9 70
69
11,193
FPL’23 (Que et al., 2023)
MetaML
- VU9P 76.1 70
638
69,751
FPL’23 (Que et al., 2023)
MetaML
{0.05, 2, 1} VU9P 75.6 45
50
6,698
FPL’23 (Que et al., 2023)
MetaML
{0.05, 2, 4} VU9P 72.8 40
23
7,224
This work
(Best Acc.)
{0.5, 0.1, 0.1} VU9P 76.2 55
272
14,580
This work
(Best DSP)
{0.5, 3, 4} VU9P 69.7 40
5
9,026
This work
(Best LUT)
{2, 5, 1} VU9P 72.1 45
47
4,410
This workb
(Acc.-DSP-LUT)
{0.5, 2, 0.5} VU9P 76.1 50
70
13,042
This workb
(Acc.-DSP-LUT)
{4, 0.5, 0.5} VU9P 75.7 45
50
6,634
This workc
(Acc.-LUT)
{4, 4, 0.5} VU9P 73.7 45
54
5,624
This workd
(Acc.-DSP)
{2, 3, 4} VU9P 70.9 40
12
7,637
@@ 5.10. Discussion and Comparison
Compared to original Jet-DNN (Duarte et al., 2018),, our design achieves higher accuracy (up to 76.2%) while using less hardware resources. When compared with LogicNets JSC-M and JSC-L (Umuroglu et al., 2020), which achieve accuracies of 70.6% and 71.8% respectively, our design demonstrates up to 5.6% higher accuracy while also offering resource efficiency. Against the AutoQkeras Q6 and QE designs (Coelho et al., 2021), which yield accuracies of 74.8% and 72.3% respectively, our framework attains higher accuracy and lower latency while providing more granular trade-offs between resource utilization and latency. Compared to MetaML (Que et al., 2023) with manual optimization, our work achieves comparable or better accuracy across various configurations with automatic optimization. Overall, the versatility of our design, which balances accuracy, latency, and resource utilization across different objectives, highlights its superiority and adaptability for diverse FPGA-based DNN applications.
@@ 6. Conclusion
This paper presents a novel co-optimization framework for FPGA-based DNN accelerators, which comprises building blocks that facilitate rapid development of customized cross-stage design flows, automating the entire design iteration process. The results demonstrate that our approach significantly reduces DSP resource usage by up to 92% and LUT usage by up to 89%, while maintaining accuracy, without requiring human effort or domain expertise. In addition, our results reveal that Bayesian optimization in DSE significantly streamlines the process, achieving results comparable to grid search but with a 15.6-fold reduction in processing time. Our futurework will extend this approach to include RTL space, support more DNN architectures, including Variational Autoencoder (VAE) (Que et al., 2024b) and transformer (Wojcicki et al., 2022) networks, and incorporate more optimization strategies like balancing initiation interval (Que et al., 2021; Rognlien et al., 2022) for further improvements in hardware efficiency and model performance.
Acknowledgement: The support of the United Kingdom EPSRC (grant numbers EP/V028251/1, EP/L016796/1, EP/N031768/1, EP/P010040/1, EP/S030069/1, and EP/X036006/1), Intel, and AMD is gratefully acknowledged. We thank Markus Rognlien, Shuo Liu and Anyan Zhao for their help.

### v2

@@ 5.2. Single $O$-task Strategies
VGG7 Baseline - 4568 (37.2%) 345k (20.0%) 65k (1.9%) 98.2%
VGG7 Quant. 0.01 934 (7.6%) 313k (18.1%) 51k (1.5%) 97.5%
VGG7 Quant. 0.05 505 (4.1%) 318k (18.4%) 66k (1.9%) 93.9%
JetDNN Baseline (Que et al., 2023) - 638 (290%) 66k (124%) 30k (28%) 76.1%
JetDNN Quant. 0.01 75 (34.1%) 57k (107%) 17k (16%) 75.5%
JetDNN Quant. 0.05 61 (27.7%) 76k (144%) 7.2k (6.8%) 71.6%
@@
5.3. Multiple $O$-task Strategies
@@ 5.6. Bottom-up Flow
5.7. Combination and Order
@@
5.8. Tolerable Loss Variation
@@
Model $\alpha_{s}$, $\alpha_{p}$, $\alpha_{q}$ (%) FPGA Acc.(%) Lat. (ns) DSP (%) LUT (%)
HLS4ML Jet-DNN (Duarte et al., 2018) - KU115 75 75 954 (17.3) -
FPL’20 LogicNets JSC-M (Umuroglu et al., 2020) - VU9P 70.6 NA 0 (0) 14,428 (1.2)
FPL’20 LogicNets JSC-L (Umuroglu et al., 2020) - VU9P 71.8 13a 0 (0) 37,931 (3.2)
NMI’21 (Coelho et al., 2021) Qkeras Q6 - VU9P 74.8 55 124 (1.8) 39,782 (3.4)
NMI’21 (Coelho et al., 2021) AutoQkeras QE - VU9P 72.3 55 66 (1.0) 9,149 (0.8)
NMI’21 (Coelho et al., 2021) AutoQkeras QB - VU9P 71.9 70 69 (1.0) 11,193 (0.9)
PL’23 (Que et al., 2023) MetaML - VU9P 76.1 70 638 (9.3) 69,751 (5.9)
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 1} VU9P 75.6 45 50 (0.7) 6,698 (0.6)
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 4} VU9P 72.8 40 23 (0.2) 7,224 (0.6)
This work (Best Acc.) {0.5, 0.1, 0.1} VU9P 76.2 55 272 (4.0) 14,580 (1.2)
This work (Best DSP) {0.5, 3, 4} VU9P 69.7 40 5 (0.1) 9,026 (0.8)
This work (Best LUT) {2, 5, 1} VU9P 72.1 45 47 (0.7) 4,410 (0.4)
This workb (Acc.-DSP-LUT) {0.5, 2, 0.5} VU9P 76.1 50 70 (1.0) 13,042 (1.1)
This workb (Acc.-DSP-LUT) {4, 0.5, 0.5} VU9P 75.7 45 50 (0.7) 6,634 (0.6)
This workc (Acc.-LUT) {4, 4, 0.5} VU9P 73.7 45 54 (0.8) 5,624 (0.5)
This workd (Acc.-DSP) {2, 3, 4} VU9P 70.9 40 12 (0.2) 7,637 (0.8)
@@ 5.10. Discussion and Comparison
Compared to original Jet-DNN (Duarte et al., 2018), our design achieves higher accuracy (up to 76.2%) while using less hardware resources. When compared with LogicNets JSC-M and JSC-L (Umuroglu et al., 2020), which achieve accuracies of 70.6% and 71.8% respectively, our design demonstrates up to 5.6% higher accuracy while also offering resource efficiency. Against the AutoQkeras Q6 and QE designs (Coelho et al., 2021), which yield accuracies of 74.8% and 72.3% respectively, our framework attains higher accuracy and lower latency while providing more granular trade-offs between resource utilization and latency. Compared to MetaML (Que et al., 2023) with manual optimization, our work achieves comparable or better accuracy across various configurations with automatic optimization. Overall, the versatility of our design, which balances accuracy, latency, and resource utilization across different objectives, highlights its superiority and adaptability for diverse FPGA-based DNN applications.
@@ 6. Conclusion
This paper presents a novel co-optimization framework for FPGA-based DNN accelerators, which comprises building blocks that facilitate rapid development of customized cross-stage design flows, automating the entire design iteration process. The results demonstrate that our approach significantly reduces DSP resource usage by up to 92% and LUT usage by up to 89%, while maintaining accuracy and without requiring human effort or domain expertise. In addition, our results reveal that Bayesian optimization in DSE significantly streamlines the process, achieving results comparable to grid search but with a 15.6-fold reduction in processing time. Future work will extend this approach to include RTL space, support more DNN architectures, including Variational Autoencoder (VAE) (Que et al., 2024b) and transformer (Wojcicki et al., 2022) networks, and incorporate more optimization strategies like balancing initiation interval (Que et al., 2021; Rognlien et al., 2022) for further improvements in hardware efficiency and model performance.
Acknowledgement: The support of the United Kingdom EPSRC (grant number UKRI256, EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1), Intel, and AMD is gratefully acknowledged. We thank Markus Rognlien, Shuo Liu and Anyan Zhao for their help.